# Transferable Mask Transformer: Cross-domain Semantic Segmentation with Region-adaptive Transferability Estimation

### v2

@@ Abstract
Recent advances in Vision Transformers (ViTs) have set new benchmarks in semantic segmentation. However, when adapting pretrained ViTs to new target domains, significant performance degradation often occurs due to distribution shifts, resulting in suboptimal global attention. Since self-attention mechanisms are inherently data-driven, they may fail to effectively attend to key objects when source and target domains exhibit differences in texture, scale, or object co-occurrence patterns. While global and patch-level domain adaptation methods provide partial solutions, region-level adaptation with dynamically shaped regions is crucial due to spatial heterogeneity in transferability across different areas of the image. In this paper, we present Transferable Mask Transformer (TMT), a novel region-level adaptation framework for semantic segmentation, which aligns cross-domain representations through spatial transferability analysis. TMT consists of two key components: (1) An Adaptive Cluster-based Transferability Estimator (ACTE), which dynamically segments images into structurally and semantically coherent regions for localized transferability assessment, and (2) A Transferable Masked Attention (TMA) module, which integrates region-specific transferability maps into ViTs’ attention mechanisms, prioritizing adaptation in regions with low transferability and high semantic uncertainty. Comprehensive evaluations across 20 cross-domain pairs demonstrate TMT’s superiority, achieving an average $2\%$ MIoU improvement over vanilla fine-tuning and a $1.28\%$ increase compared to the state-of-the-art baselines. The source code will be publicly available soon.
@@ 1 Introduction
Figure 1: Illustration of Different Datasets and Attention Mechanisms: Fig.a shows the domain gap across different datasets, highlighting the challenges encountered during data collection in various real-world scenarios. In Fig.b-c, the attention mask and attention map visualizations indicate that vanilla fine-tuning models tend to predict misleading masks, focusing on less relevant regions (highlighted in red). In contrast, our method generates more accurate masks (shown in blue), directing attention to more critical regions in the target task, such as the car. Fig.d compares two region partition methods for estimating transferability, where our approach segments the scene into adaptive regions.
@@
While existing methods for leveraging transferability have demonstrated success in CNN-based architectures [23], their direct application to vision transformers remains underexplored. To bridge this gap, we propose Transferable Masked Attention (TMA), a novel module tailored for transformer-based models. TMA dynamically adjusts attention masks by incorporating region-specific transferability maps, enabling object queries to focus on regions with both low transferability and low semantic confidence. By combining transferability scores in a hierarchical way, TMA ensures that the model prioritizes regions requiring adaptation while suppressing attention to well-aligned or confidently predicted areas. This approach not only mitigates inappropriate mask predictions but also enhances the model’s ability to capture domain-specific details, offering a principled way to integrate transferability estimation into the attention mechanism of vision transformers.
@@
Our experimental results show that the proposed methods achieve an average MIoU improvement of 2% compared to vanilla fine-tuning and 1.28% over the SOTA method across 20 source-target transfer pairs over 5 popular semantic segmentation benchmarks for autonomous driving.
Figure 2: Overview of the framework. The model training begins with the ACTE, which is first trained using both source and target data. The lower section illustrates how ACTE evaluates and assigns different transferability scores to various regions, represented by different colors in the transferability map. After ACTE has been trained, its output—region-level transferability maps—is used to guide the training of the TMA within attention mechanism (top-right).
2 Related Work
2.1 Transformer-Based Visual Segmentation
Visual segmentation is a fundamental problem in computer vision and involves numerous real-world applications, such as robotics, social media, autonomous driving, etc. [18]. Recently, researchers applied transformers to Computer Vision (CV) tasks. Vision Transforemer (ViT) directly takes the sequence of image patches to classify the full image [7]. It achieves state-of-the-art performance on multiple image recognition datasets. DETR [2] simplified object detection with object queries. Mask2Former [5] further improved training efficiency by introducing mask attention, reducing query ambiguity, and limiting cross-attention scope. Despite the success of Mask2Former, challenges arise in domain adaptation, where attention masks derived from the source domain may not generalize well to different target domains due to variations in visual features.
2.2 Transfer Learning in Transformers
Before ViTs, domain-invariant representation learning strategies were used [32]. DAFormer introduced a transformer backbone with rare class sampling, ImageNet feature loss, and learning rate warm-up, becoming a strong baseline [11]. HRDA [12] improved on DAFormer with multi-resolution training, utilizing different crops to maintain fine segmentation details. MIC [13] applied masked image consistency to enhance spatial context relations by enforcing consistency between masked target images and pseudo-labels through a teacher-student framework. SFA [27] and DADETR [34] focused on feature alignment with transformers, employing domain-specific queries and hybrid attention modules. These methods are generally based on unsupervised domain adaptation strategies, aiming to reduce distribution discrepancies between the source and target domains.
2.3 Transferability Estimation in Vision Tasks
Recent advancements in transferability estimation have significantly improved the efficiency and accuracy of transfer learning in vision tasks. OTCE metrics [24] introduced auxiliary-free methods to evaluate transferability with reduced computational costs. Yang et al. extended these concepts to semantic segmentation by using pixel-wise transferability scores, improving fine-tuning accuracy by focusing on challenging low-transferability regions [23]. In object detection, Chen et al. propose HTCN [3] to balance transferability and discriminability, optimizing feature representations at multiple levels. Yang et al. introduced TADA [31], a method that aligns only the most transferable regions within images. Lastly, Yang et al. proposed the Transferable Vision Transformer (TVT), leveraging ViTs to enhance transferability in unsupervised domain adaptation, which differs from ours as we employ supervised fine-tuning.
3 Preliminary
3.1 Supervised Finetuning
For segmentation tasks, the input space $X=\{\bm{x}\colon P\to\mathbb{R}\}$ is the image set, and the output space $Y=\{\bm{y}\colon P\to\Lambda\}$ is the set of partition of images in label, $P=\{1,\dots,h\}\times\{1,\dots,w\}$ is the pixel set. Source domain $\mathcal{D}_{\mathrm{S}}$ and target domain $\mathcal{D}_{\mathrm{T}}$ are different distributions over $X\times Y$. Given a source dataset $\mathcal{S}=\{(\bm{x}_{\mathrm{S}}^{i},\bm{y}_{\mathrm{S}}^{i})\}_{i=1}^{n}$ drawn i.i.d. from $\mathcal{D}_{\mathrm{S}}$ and a target dataset $\mathcal{T}=\{(\bm{x}_{\mathrm{T}}^{i},\bm{y}_{\mathrm{T}}^{i})\}_{i=1}^{n^{\prime}}$ drawn i.i.d. from $\mathcal{D}_{\mathrm{T}}$. We aim to build a classifier $\sigma\colon X\to Y$ with a low target risk
$R_{\mathcal{D}_{\mathrm{T}}}(\sigma):=\operatorname{E}\mathop{\lvert}\{p\in P\mathop{|}\sigma(\bm{x}_{\mathrm{T}})(p)\neq\bm{y}_{\mathrm{T}}(p)\}\mathop{\rvert}$, (1)
in which $(\bm{x}_{\mathrm{S}},\bm{y}_{\mathrm{S}})\sim\mathcal{D}_{\mathrm{S}},\ (\bm{x}_{\mathrm{T}},\bm{y}_{\mathrm{T}})\sim\mathcal{D}_{\mathrm{T}}$. Fine-tuning involves initializing the target model $\theta_{t}$ with the source model parameters $\theta_{s}$, followed by supervised learning on the target data. In the context of supervised finetuning for segmentation tasks, the target risk $R_{\mathcal{D}_{\mathrm{T}}}(\sigma)$ quantifies the expected prediction error of the classifier $\sigma$ on the target domain. Specifically, it measures the average number of misclassified pixels across the target dataset. This risk is typically computed using a pixel-wise loss function, such as the Cross-Entropy Loss or Dice Loss.
3.2 Measuring Transferability with Proxy A-distance
Let $\mathcal{D}_{{}^{X}}$ denote the marginal distribution of $\mathcal{D}$ over $X$. Let $\mathcal{D}_{{}^{X}}$ denote the marginal distribution of $\mathcal{D}$ over $X$. The transferability between domains $\mathcal{D}_{\mathrm{S}}^{X}$ and $\mathcal{D}_{\mathrm{T}}^{X}$ is measured using Proxy $\mathcal{A}$-distance (PAD) [8]. This distance metric quantifies the divergence between source and target domains, providing a principled approach to assess domain shift.
Given $\mathcal{A}$ as a collection of measurable sets on $X$, the $\mathcal{A}$-divergence between $\mathcal{D}_{\mathrm{S}}^{X}$ and $\mathcal{D}_{\mathrm{T}}^{X}$ is given by
@@ 3.2 Measuring Transferability with Proxy A-distance
We estimate it with the $\mathcal{A}$-divergence between $\mathcal{S}^{X}$ and $\mathcal{T}^{X}$,
$\begin{split}&\!\!\!\!d_{\mathcal{A}}(\mathcal{S}^{X},\mathcal{T}^{X})=\\ &\!\!\!\!2\Bigl{[}1-\min_{A\in\mathcal{A}}\Bigl{(}\frac{1}{n}\!\sum_{1\leqslant i\leqslant n}\!\mathrm{I}[\bm{x}_{\mathrm{S}}^{i}\!\notin\!A]+\frac{1}{n^{\prime}}\!\sum_{1\leqslant i\leqslant n^{\prime}}\!\mathrm{I}[\bm{x}_{\mathrm{T}}^{i}\!\in\!A]\Bigr{)}\Bigr{]},\end{split}$ (3)
in which $\mathcal{S}^{X}=\{\bm{x}\mathop{|}\mathop{\exists}\bm{y}\operatorname{s.t.}(\bm{x},\bm{y})\in\mathcal{S}\}$ for a dataset $\mathcal{S}$. $\mathrm{I}[p]$ is the indicator function. If predicate $p$ is true, $\mathrm{I}[p]=1$, otherwise $\mathrm{I}[p]=0$.
Although computing $d_{\mathcal{A}}(\mathcal{S}^{X},\mathcal{T}^{X})$ exactly is generally challenging, an approximation can be efficiently obtained by training a classifier to distinguish between source and target examples [15]. Define the labeled dataset as
$\mathcal{L}=\{(\bm{x}_{\mathrm{S}}^{i},0)\}_{i=1}^{n}\cup\{(\bm{x}_{\mathrm{T}}^{i},1)\}_{i=1}^{n^{\prime}},$ (4)
where source and target samples are labeled 0 and 1, respectively. This yields the Proxy $\mathcal{A}$-distance as an estimate of $\mathcal{A}$-divergence:
$\hat{d}_{\mathcal{A}}=2(1-2\epsilon).$ (5)
A domain discriminator $E$ trained on $\mathcal{L}$ provides a generalization error $\epsilon$, which approximates the minimization term in Equation (3).
4 Method
Our framework, TMT, illustrated in Fig. 2, builds upon the Mask2Former architecture to address cross-domain image segmentation through a unified and layer-wise adaptive pipeline. The process begins with feature extraction, where a backbone network captures hierarchical features from the input image, followed by a pixel decoder that refines and upsamples these features to preserve spatial details. These feature maps are then processed by the Adaptive Clustering and Transferability Estimator (ACTE), which adaptively segments the image into regions based on structural characteristics and assigns transferability scores to each region. These scores, encoded in a transferability map, quantify domain alignment and guide the subsequent learning process. Crucially, our Transferable Masked Attention (TMA) mechanism integrates this transferability map into every layer of the network, dynamically modulating the attention mechanism to enhance domain adaptation at multiple feature levels. Finally, the refined features and masks are used to generate class predictions and segmentation outputs, enabling accurate and domain-adaptive image segmentation. By incorporating transferability-aware modulation at each layer, our end-to-end pipeline ensures robust performance across domains, explicitly addressing domain shifts at both regional and hierarchical levels.
4.1 Adaptive Cluster Transferability Estimator
Adaptive Clustering Method.  To capture inter-regional structural differences while preserving intra-regional structural and semantic consistency, our approach generates adaptively sized regions that align closely with image boundaries. As illustrated in Fig.3, we propose an adaptive clustering method inspired by the classic works on superpixel learning [1, 17], which significantly reduces the time complexity by confining the computation range of the pixel similarity matrix. We perform an iterative grouping algorithm that takes a feature map $\bm{\chi}$ as input, and gives out a region-level feature map $\bm{f}$, where $R$ denotes the set of regions. We define a soft assignment $\text{\large$a$}\colon P\times R\to[0,1]$, where $\text{\large$a$}(p,r)$ indicates the probability of pixel $p$ being assigned to region $r$.
Figure 3: Pixel features are compared to nearby center features to compute similarities. Each pixel is assigned to one of its nearby center features based on their similarities.
Initially, the pixel space $P$ is divided into cells of size $c\times c$, which serve as clustering centers. To assign coordinates to regions, the region set is defined as $R=\{1,\dots,h/c\}\times\{1,\dots,w/c\}$. Each pixel is assigned to its corresponding cell, resulting in the initial assignment $a$ as follows:
$\text{\large$a$}((x,\!y),(i,\!j))=\mathrm{I}[\text{$(i\!-\!1)c\hskip-0.83328pt<\hskip-0.83328ptx\hskip-0.83328pt\leqslant\hskip-0.83328ptic\text{ and }(j\!-\!1)r\hskip-0.83328pt<\hskip-0.83328pty\hskip-0.83328pt\leqslant\hskip-0.83328ptjr$}].$ (6)
The algorithm then repeats following steps for $L$ iterations. First, the region-level feature map $\bm{f}$ is computed by averaging the features within each region:
$\bm{f}(r)=\frac{1}{\sum_{p\in P}\text{\large$a$}(p,r)}\sum\nolimits_{p\in P}\text{\large$a$}(p,r)\bm{\chi}(p).$ (7)
Next, the similarity between pixels and clustering centers is computed. For computational efficiency, we limit this comparison to the $3\times 3$ grid of nearby centers surrounding each pixel, as illustrated in Fig.3. The similarity function $s\colon P\times R\to[0,+\infty)$ is defined as:
[…]
Here, $\kappa$ is a hyperparameter that controls the sharpness of the similarity distribution. This approach ensures that clustering centers only aggregate nearby pixels, forming cohesive region-level image segments. The new soft assignment is then computed as:
$\text{\large$a$}_{\text{new}}(p,r)=\frac{s(p,r)}{\sum_{\pi\in R}s(p,\pi)}.$ (9)
This updated assignment is substituted back into Equation (7) to recalculate the region-level feature map $\bm{f}$. After $\bm{L}$ iterations, the process converges to a final cluster region feature map. Each position’s feature vector $\bm{f}(r)$ corresponding to a region with a certain receptive field in the original image.
Region-level Transferability Estimator.  Building upon the theoretical framework of Proxy A-distance, we propose a region-level transferability estimation mechanism. To extend this concept to region-level adaptation, we design an domain discriminator $E$ that operates on adaptive regions $\bm{f}(r)$ rather than entire images. Given region-level features $\bm{f}(r)$ extracted from both domains, $E$ predicts the probability $E(\bm{f}(r))\in[0,1]$ that region $r$ originates from the target domain. To calculate the generalization error $\epsilon$ in Equation (5), we can use the cross-entropy loss, defined by:
$\mathcal{L}(E(\bm{f}(r)),d)=(1-d)\log\frac{1}{1\!-\!E(\bm{f}(r))}+d\log\frac{1}{E(\bm{f}(r))},$ (10)
where $d$ represents the domain label of image $\bm{x}\in\mathcal{L}^{X}$ that generates $\bm{f}$, indicating whether $\bm{x}$ is drawn from the source domain ($d=0$) or from the target domain ($d=1$). To train $E$, we construct a training set consisting of features $\bm{F}$ extracted from both the source and target domains, with corresponding domain labels $d$. Training data is constructed by sampling balanced batches from both domains, with features $\bm{f}(r)$ generated by our adaptive cluster module.
A lower value of $E(\bm{f}(r))$ indicates that the features are more similar to the source domain, which is labeled as 0. So we define the transferability map $\bm{T}$ for each pixel as:
$\bm{T}_{x,y}=1-E\left(\bm{f}\left(r^{*}\right)\right),$ (11)
where $r^{*}=\operatorname*{arg\ max}_{r\in R}\text{\large$a$}((x,y),r)$. A high $\bm{T}_{x,y}$ value indicates near-perfect domain alignment, suggesting that the region’s feature distribution closely matches the source domain. Such regions require minimal adaptation, as they already exhibit strong transferability and provide limited room for improvement through fine-tuning. Conversely, regions with lower $\bm{T}_{x,y}$ exhibit significant distribution shifts from the source domain, necessitating substantial and targeted fine-tuning to bridge the domain gap. By focusing on these low-transferability regions, we can effectively adapt the model to capture the unique characteristics of the target domain, thereby enhancing the model’s cross-domain generalization ability.
@@ 4.2 Transferability Guided Masked Attention
  Domain MIoU fwIoU mACC
[…]
Table 1: Performance comparison of different methods across various datasets when transferred from different source domains.

### v3

# TMT: Cross-domain Semantic Segmentation with Region-adaptive Transferability Estimation

@@ Abstract
Recent advances in Vision Transformers (ViTs) have significantly advanced semantic segmentation performance. However, their adaptation to new target domains remains challenged by distribution shifts, which often disrupt global attention mechanisms. While existing global and patch-level adaptation methods offer some improvements, they overlook the spatially varying transferability inherent in different image regions. To address this, we propose the Transferable Mask Transformer (TMT), a region-adaptive framework designed to enhance cross-domain representation learning through transferability guidance. First, we dynamically partition the image into coherent regions—grouped by structural and semantic similarity—and estimates their domain transferability at a localized level. Then, we incorporate region-level transferability maps directly into the self-attention mechanism of ViTs, allowing the model to adaptively focus attention on areas with lower transferability and higher semantic uncertainty. Extensive experiments across 20 diverse cross-domain settings demonstrate that TMT not only mitigates the performance degradation typically associated with domain shift but also consistently outperforms existing approaches.
Index Terms—  Cross-Domain Adaptation, Transfer Learning, Transferability Estimation, Semantic Segmentation
@@ 1 Introduction
Fig. 1: Fig.a illustrates domain gaps across datasets, reflecting real-world challenges. Fig.b–c show that vanilla fine-tuning produces misleading attention masks (red), focusing on irrelevant areas, while our method yields accurate masks (blue) by emphasizing task-critical regions like the car. Fig.d compares region partition methods, with our approach adaptively segmenting images for better transferability estimation.
Fig. 2: Overview of the framework. The model training begins with the ACTE, which is first trained using both source and target data. The lower section illustrates how ACTE evaluates and assigns different transferability scores to various regions, represented by different colors in the transferability map. After ACTE has been trained, its output—region-level transferability maps—is used to guide the training of the TMA within attention mechanism (top-right).
@@
While existing methods for leveraging transferability have demonstrated success in CNN-based architectures [13], their direct application to vision transformers remains underexplored. To bridge this gap, we propose Transferable Masked Attention (TMA) tailored for transformer-based models. TMA dynamically adjusts attention masks by incorporating region-specific transferability maps, enabling object queries to focus on regions with both low transferability and low semantic confidence. By combining transferability scores in a hierarchical way, TMA ensures that the model prioritizes regions requiring adaptation while suppressing attention to well-aligned or confidently predicted areas. This approach not only mitigates inappropriate mask predictions but also enhances the model’s ability to capture domain-specific details, offering a principled way to integrate transferability estimation into the attention mechanism of vision transformers.
@@
Our experimental results show that the proposed methods achieve an average MIoU improvement of 2.1% compared to vanilla fine-tuning and 1.28% over the SOTA method across 20 source-target transfer pairs over 5 popular semantic segmentation benchmarks for autonomous driving.
2 Method
Our framework, TMT, illustrated in Fig. 2, builds upon the Mask2Former architecture to address cross-domain image segmentation through a unified and layer-wise adaptive pipeline. After extracting hierarchical feature maps, the model adaptively segments the image into regions based on structural characteristics. A Transferability Estimator is then employed to assess region-wise transferability. This transferability map is integrated into every layer of the network, dynamically modulating the attention mechanism to enhance domain adaptation across multiple feature levels.
2.1 Adaptive Cluster
Our adaptive part-level grouping method generates structurally coherent regions aligned with image boundaries while preserving semantic consistency within each region. Following the common practice in superpixel learning, we employ an iterative grouping algorithm that processes feature maps to produce part-level mask features. The pixel space is initially divided into a regular grid of size $r\times r$, where each cell serves as an initial cluster center, yielding $N_{p}=\frac{H}{r}\times\frac{W}{r}$ centers. During each iteration, we first compute the cluster center features $\mathbf{Q}\in\mathbb{R}^{N_{p}\times d}$ by averaging the pixel features within each region. Next, we compute the cosine similarity between each pixel and its surrounding $3\times 3$ centers. The similarity between the $i$-th center feature $\mathbf{Q}_{i}$ and the $j$-th pixel feature $\mathbf{K}_{j}$ is given by:
$D_{i,j}=\begin{cases}f(\mathbf{Q}_{i},\mathbf{K}_{j})=\dfrac{1}{\tau}\cdot\dfrac{\mathbf{Q}_{i}\cdot\mathbf{K}_{j}}{\|\mathbf{Q}_{i}\|\|\mathbf{K}_{j}\|}&\text{if }i\in\mathcal{N}_{j}\\ -\infty&\text{if }i\notin\mathcal{N}_{j}\end{cases}$ (1)
where $\mathcal{N}_{j}$ denotes the set of nearby centers of the $j$-th pixel, and $\tau$ is a temperature parameter.
The soft assignment matrix $\mathbf{A}\in[0,1]^{N_{p}\times(HW)}$ is then derived by applying softmax normalization over the similarity values for each pixel:
$\mathbf{A}_{i,j}=\frac{\exp(D_{i,j})}{\sum_{i=1}^{N_{p}}\exp(D_{i,j})}$ (2)
The cluster center features are updated via:
$\mathbf{Q}^{\text{new}}=\mathbf{A}\times\mathbf{K}$ (3)
This iterative refinement process converges to a final clustered feature map where each region corresponds to a coherent image segment, effectively reducing computational complexity through local constraints while maintaining structural alignment with image content.
2.2 Region-level Transferability Estimator
To quantify the domain shift between source and target domains, we adopt the Proxy $\mathcal{A}$-distance (PAD) [14], which measures the divergence between the marginal distributions $\mathcal{D}_{\mathrm{S}}^{X}$ and $\mathcal{D}_{\mathrm{T}}^{X}$. The $\mathcal{A}$-divergence is defined as:
@@ 2.2 Region-level Transferability Estimator
In practice, we estimate this divergence by training a domain discriminator $E$ to classify samples from the source and target domains. The Proxy $\mathcal{A}$-distance is then approximated based on the generalization error of this discriminator. We extend this PAD framework to operate at the region level rather than at the image level, enabling a fine-grained assessment of transferability across different regions within each image.
We design a domain discriminator $E$ that takes region-level features $\mathbf{Q}_{i}$ from both domains as input and predicts the probability that the region originates from the target domain. The discriminator is optimized using a cross-entropy loss:
$\mathcal{L}(E(\mathbf{Q}_{i}),d)=(1-d)\log\frac{1}{1-E(\mathbf{Q}_{i})}+d\log\frac{1}{E(\mathbf{Q}_{i})},$ (5)
where $d$ denotes the domain label, with $d=1$ for source domains and $d=0$ for the target domain. To train $E$, we construct a balanced training set comprising region features from both domains, each annotated with their corresponding domain labels. Batches are uniformly sampled from both domains to ensure balanced supervision during training.
Regions with high $\bm{T}_{i}$ require little adaptation, as they already exhibit strong transferability and offer limited potential for further improvement. In contrast, regions with lower $\bm{T}_{i}$ exhibit substantial distribution shifts relative to the source domain and therefore require more targeted adaptation to reduce the domain gap. By focusing on these low-transferability regions, we can more effectively adapt the model to the target domain’s characteristics.
@@ 2.3 Transferability Guided Masked Attention
  Method C B M S G Avg
B M S G C M S G C B S G C B M G C B M S
  MIoU Linear 55.0 67.3 43.8 56.7 47.3 47.0 29.3 39.8 72.4 61.9 48.8 64.4 31.1 22.7 28.7 28.4 38.3 30.2 40.9 27.8 44.1
DAF 57.8 67.8 52.5 59.9 50.0 46.9 36.2 41.1 72.8 62.1 53.6 65.2 36.0 26.9 34.4 33.4 41.8 32.7 43.6 39.5 47.7
MIC 60.6 69.8 66.3 66.3 53.1 47.0 47.3 47.8 73.0 62.6 60.1 70.6 45.4 31.7 38.2 43.1 51.6 38.7 45.9 48.0 53.4
Full 62.2 70.7 75.8 69.2 54.7 46.9 53.0 52.0 73.6 61.2 67.3 71.7 50.9 37.6 43.7 49.7 57.0 42.3 48.4 57.2 57.3
LEEP 62.8 71.1 76.2 69.5 55.2 47.3 53.4 52.5 72.6 62.7 68.5 72.5 51.2 38.0 44.0 50.0 57.5 42.8 48.8 57.6 57.7
OTCE 63.0 71.4 76.5 69.3 55.5 47.6 53.7 52.8 72.8 62.3 69.8 71.8 51.5 38.2 44.3 50.3 57.8 43.0 49.0 57.9 57.9
Ours 64.5 72.7 77.8 71.2 57.5 49.0 55.8 54.8 73.0 62.9 69.5 73.5 53.5 41.1 44.0 51.8 59.5 44.8 51.0 59.5 59.4
  mACC Linear 68.4 79.9 64.1 70.4 58.4 58.1 36.6 50.7 83.9 73.5 68.2 78.1 38.9 28.3 36.2 39.5 47.6 36.5 51.5 40.1 55.4
DAF 69.6 80.2 67.5 71.6 59.7 57.2 39.5 53.8 84.2 73.2 74.3 78.9 45.7 31.6 39.7 46.0 52.8 42.0 53.4 49.3 58.5
MIC 73.1 80.5 80.8 78.3 64.4 57.9 58.9 58.2 84.6 73.5 81.4 80.6 55.2 39.8 49.6 54.4 60.4 47.7 58.6 58.5 64.8
Full 73.8 80.7 85.2 79.9 67.5 57.1 63.4 61.8 84.9 73.1 84.9 81.7 62.5 46.0 54.0 60.2 68.9 51.2 59.7 67.7 68.2
LEEP 74.2 81.1 85.5 80.2 67.9 57.5 63.8 62.2 84.6 72.8 85.9 82.5 62.9 46.4 54.3 60.5 69.2 51.7 60.0 68.0 68.6
OTCE 74.5 81.4 85.8 80.5 68.2 57.8 64.0 62.5 84.3 72.5 87.3 83.9 63.2 46.7 54.6 60.8 69.5 52.0 60.3 68.1 68.9
Ours 75.8 82.8 87.1 81.7 70.2 59.8 66.4 64.3 85.2 73.7 87.0 83.7 65.7 48.5 56.1 62.2 71.0 54.0 62.0 69.8 70.4
Table 1: Performance comparison (MIoU and mACC) of different methods across various domain adaptation scenarios.